OpenStudy (anonymous):

Use the quadratic formula to find the roots of the equation. x2 + 3x - 18 = 0 A. {-6, 3} B. {-2, 9} C. {-3, 6} D. {-9, 2}

OpenStudy (anonymous):

( x +6) ( x -3) x= -6,3

OpenStudy (anonymous):

a-1,b=3 ,c= -18 b^2-4ac= 3^2- 4 ( 1) (-18)= 81 sq 81=9 -3+ 9/ 2=3 -3-9/2= -6